OpenStudy (anonymous):

if log36.2=1.5587 ,what is the characteristics of log=36.2

OpenStudy (anonymous):

sir.. please help me

OpenStudy (anonymous):

is the answer 1? how to get this? what's the solution sir?

OpenStudy (calculusfunctions):

Every common logarithm has two main parts: i). an integer part called the "characteristic", and ii). a decimal part called the "mantissa". The characteristic of every common logarithm depends on the argument number N. If N > 1, the characteristic is one less than the number of digits in the whole number portion of the argument number N. If 0 < N < 1, the characteristic is negative and one greater than the number of leading zeros in the decimal place of the argument N. For example, Suppose N = 2 536.002 (a number greater than 1). Then log (2 536.002) = log (1000 x 2.536 002) log (2 536.002) = log (1000) + log (2.536 002) [∵ log (xy) = log x + log y]. Continuing log (1000) + log (2.536 002) = 3 + log (2.536 002) ∴ the characteristic of log (2 536.002) is 3. Now suppose N is a number between 0 and 1. Let's say N = 0.011 235. log (0.011 235) = log (1.112 35 x 0.01) log (0.011 235) = log (1.112 35) + log (0.01) log (0.011 235) = log (1.112 35) + (-2) ∴ the characteristic of log (0.011 235) is -2 [∵ 10ˆ-2 = 0.01]. Now that we've built character, let's find the mantissa. This one's too easy. Let's take a look at the above examples one more time. Recall log (2 536.002) = 3 + log (2.536 002) ∴ log (2.536 002) = 0.404 149 591 713 ... is the mantissa @02231985 as a teacher I won't simply give you the answer. I will teach you and help you learn. I wrote the short lesson above for you, now you read it and try it yourself.

OpenStudy (anonymous):

sir? what's the N here? the 36.2 or the1.5587

OpenStudy (anonymous):

i think 36.2

OpenStudy (calculusfunctions):

Read the lesson carefully and you shall see that the N is the argument or the antilogarithm. Thus in log (36.2), N = 36.2.

OpenStudy (anonymous):

yeah i'm right

OpenStudy (calculusfunctions):

Yes you are smart!!!!!!! :)

OpenStudy (anonymous):

log (1000 ) but i'm confused. . where did you get that? is that from \[\log {2536.002} =1000\]

OpenStudy (anonymous):

or is it constsant for numbers greater than1

OpenStudy (calculusfunctions):

Because log (2536.002) = log (1000 X 2.536.002)

OpenStudy (anonymous):

yeah i saw its like it is expanded of the simpler theory of log (xy) = log x + log y but sir. . the x there is 1000 is 1000 constant in determining characteristics of numbers greater than 1

OpenStudy (calculusfunctions):

Let me explain again, but in different terms. To find the mantissa, write the argument in scientific notation form, then the mantissa will be the power of 10. Thus if given log (2536.002), the argument is 2536.002. Now written in scientific notation, 2536.002 = 2.536 002 X 10³ Here the power of 10 is 3. ∴ the mantissa is 3.

OpenStudy (calculusfunctions):

Oh I see you already got it! Great! Good Work!!

OpenStudy (anonymous):

ahhhhh like in scientific notations

OpenStudy (anonymous):

my answer above is 1 is it right sir?

OpenStudy (calculusfunctions):

Sorry, I mean characteristic, NOT mantissa!

OpenStudy (anonymous):

sir? ? ? is my answer right?

OpenStudy (calculusfunctions):

Yes your answer of 1 is correct! Log (36.2) = log (10 X 3.62) log (36.2) = log 10 + log (3.62) log (36.2) = 1 + log (3.62) ∴ the characteristic is 1. OR 36.2 in scientific notation is 3.62 X 10. Since we have the first power of ten, the characteristic is 1. Good job @02231985

OpenStudy (anonymous):

sir can i ask another question?

OpenStudy (anonymous):

thank you @calculusfunctions

OpenStudy (calculusfunctions):

Welcome! Sure you may ask but I have to go soon.

OpenStudy (anonymous):

yes sir. hm The half-life of a radioactive substance is 5 minutes. how much of an original mass of 100 grams remains after half an hour?

OpenStudy (calculusfunctions):

You may use the universal half-life formula\[M =M _{0}\left( \frac{ 1 }{ 2 } \right)^{\frac{ t }{ h }}\]where\[M _{0}\] is the initial mass present,\[h\]is the time it takes for the initial mass to reach half of it,\[t\]is the time it takes the substance to decay to the final mass, and \[M\]is the final mass.

OpenStudy (calculusfunctions):

You may also apply calculus but I presume you may not have learned calculus yet.

OpenStudy (calculusfunctions):

One quick note: t and h must have the same unit of time.

OpenStudy (calculusfunctions):

@02231985 are you there? If you understand, then please reply.

OpenStudy (anonymous):

sir . . i'm confused on where I could substitute the 5 in t or h? but sir even i Substitute those two : t=30 h=5

OpenStudy (calculusfunctions):

Yes you have it right! t = 30 minutes h = 5 minutes because the question very clearly stated that the substance has a half life of 5 minutes.

OpenStudy (anonymous):

i have used that also. . . its just that \[M _{0} is N _{0}\] the answers i got was wrong it was not in the choices sir. .

OpenStudy (calculusfunctions):

So did you get it?

OpenStudy (anonymous):

\[100(\frac{ 1 }{ 2 })^{\frac{ 30 }{ 5}}\]

OpenStudy (calculusfunctions):

Yes M equals the expression you have above. So what do you get for M?

OpenStudy (anonymous):

sir should i get first the (12)305

OpenStudy (calculusfunctions):

Remember the Order of Operations (BEDMAS) Always!!

OpenStudy (anonymous):

(\[\frac{ 1 }{ 2})^{6}\]

OpenStudy (calculusfunctions):

Yes! Then multiply by 100.

OpenStudy (calculusfunctions):

So what is your final mass?

OpenStudy (anonymous):

\[1\frac{ 9 }{ 16 }\]

OpenStudy (calculusfunctions):

No. It is not a mixed fraction!

OpenStudy (anonymous):

1.5625

OpenStudy (calculusfunctions):

100 X (1/2)ˆ6

OpenStudy (anonymous):

1.5625

OpenStudy (calculusfunctions):

Yes! That is the final mass! 1.5625 grams. Very Good!

OpenStudy (anonymous):

ah yes SIR is it morning there?

OpenStudy (calculusfunctions):

Sorry, your mixed fraction was right but what I meant to say is that for this type of question, express your answer as a decimal. Good Job!

OpenStudy (anonymous):

yes sir

OpenStudy (calculusfunctions):

Yes it is! Where are you from?

OpenStudy (anonymous):

philippines sir. .

OpenStudy (calculusfunctions):

I see. I'm in Canada.

OpenStudy (anonymous):

Two sides of a triangular plot of land have lengths of 400ft and 600 ft. The measurement of the angle between those sides is 46degrees 20. Find the perimeter of the plot

OpenStudy (anonymous):

ahh ok sir. . thank you very much sir. .

OpenStudy (calculusfunctions):

I have to go now, otherwise I'd definitely help. But quickly, what do you mean by 46 degrees 20? Do you mean 46 degrees and 20 minutes? Please be clear.

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

sir can i leave message to you about those questions. . . i will just read the lesson sir. . .

OpenStudy (calculusfunctions):

Yes what? 46 degrees and 20 minutes?

OpenStudy (anonymous):

yes sir 46 degrees and 20 minutes

OpenStudy (calculusfunctions):

OK so quickly before I go I'll help you with this last one. Are you ready?

OpenStudy (anonymous):

yes sir

OpenStudy (calculusfunctions):

OK first to convert minutes into degrees, multiply the minutes by 1/60. Now draw and label the triangle using the draw feature below and show me quickly.

OpenStudy (anonymous):

1/3

OpenStudy (calculusfunctions):

20 minutes equals 20/60 or 0.333333.......... degrees. Thus 46 degrees and 20 minutes equals 46.33 degrees approximately. Understand? Now please show me your diagram.

OpenStudy (anonymous):